OpenStudy (anonymous):

2x + y = 4 3x - y = 6

OpenStudy (anonymous):

Add the equations together: 2x + 3x + y - y = 4 + 6 Simplify 5x = 10 x = 2 Now you know x, so you can solve for y by plugging x into the first equation 2(2) + y = 4 4 + y = 4 y = 0

OpenStudy (anonymous):

thank you!!